Find the diameter of an 18 m (L) long steel wire that will stretch no more than 9 mm(∆L) when a load (stress) of 3800 N is hung on the end of the wire. The Young’s Modulus for steel wire is 200 x 109 N/m2




1
Expert's answer
2022-04-07T18:35:49-0400

"\frac{\Delta L} L=\frac{F}{SE}=\frac F{\pi r^2E},"

"d=2\sqrt{\frac{FL}{\pi E\Delta L}}=7~mm."